OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution. MEDALS TO WHOEVER HAS THE BEST ANSWERS!! :)

OpenStudy (anonymous):

|2x + 8| = 6 2x+8=6 or 2x+8=-6 x=-1 or x=-7

OpenStudy (hoblos):

2x+8 =6 => x=-1 OR 2x+8 =-6 => x= -7